OpenStudy (anonymous):

Find the equation of the line that contains the given point and has the given slope. Let y be the dependent variable and let x be the independent variable. Point (4, 5), m = 2/3 this is what I figured but is not right... y-5=2/3(x-4) y=2/3x-8/3 y=2/3x-(8/3+5) y=2/3x ?

OpenStudy (anonymous):

y=2/3x-(8/3+5)

OpenStudy (anonymous):

I did do that but it is saying its wrong so I didnt know if i was doing something wrong

OpenStudy (anonymous):

do I need to reduce something

OpenStudy (anonymous):

y=2/3x-8/3 this must n=be the ans

OpenStudy (callisto):

y-5=2/3(x-4) y=2/3x-8/3 <- Wrong for this step... y-5=2/3(x-4) y-5 = (2/3)x - (2/3)(4) y-5 = (2/3)x - 8/3 Now, add 5 to both sides, can you work out the answer?

OpenStudy (anonymous):

well its 7 2/3 right?....23/3

OpenStudy (callisto):

Hmm.. you've 'accidentally' missed the -5 on the left hand side for your second step..

OpenStudy (callisto):

Yup, for the constant term only

OpenStudy (anonymous):

oh ok so with the neg it being +7/3 now for the ending

OpenStudy (callisto):

+7/3 ?

OpenStudy (anonymous):

y=2/3x+7/3

OpenStudy (anonymous):

which my assignment is now saying is correct

OpenStudy (callisto):

oop...my mistakes.. yea...you're correct!!

OpenStudy (anonymous):

ok thank you for your help

OpenStudy (callisto):

you're welcome